. A light signal is traveling through a fiber. What is the delay in the signal if the length of the fiber-optic cable is 1.5 Km (assume a propagation speed of 3 x 108 m/s)?
Answer #1

delay
= length of the cable / speed of the signal
= 1.5 km / 3*10^8 m/s
= 1.5*1000 m / 3*10^8 m/s
= 0.5 * 10^-5 seconds
= 0.005 * 10^-3 seconds
= 0.005 milli-seconds

Answer: 0.005 milli-seconds